# Global solutions to cross-diffusion systems with independent advections in one dimension

### 20 Mar 2026 v1

@@ 5.1. The support of $\pi_{t,x}$ is (almost always) at most a line
Indeed, we split $\mathbb{R}$ for [M,M][-M,M] and $\mathbb{R}\setminus[-M,M]$. On [M,M][-M,M] we have the covergence of integrals thanks to (C1), (C5) and because $\overline{\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}s_{\varepsilon}^{\alpha}}=\frac{u}{s}\,\partial_{x}s^{\alpha}$. On $\mathbb{R}\setminus[-M,M]$ we use that the integral is small
$\left|\int_{0}^{t}\int_{\mathbb{R}\setminus[-M,M]}u_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau\right|\leq\|\partial_{x}\mathcal{V}^{1}\|_{L^{\infty}}\,\|\partial_{x}s_{\varepsilon}^{\alpha}\|_{L^{2}((0,T)\times\mathbb{R})}\,\frac{\|u_{\varepsilon}\,|x|^{\kappa}\|_{L^{2}((0,T)\times\mathbb{R})}}{M^{\kappa}},$
where $\kappa=1$ when $\alpha>\frac{1}{3}$ and $\kappa=\frac{1}{2}$ when $\alpha\leq\frac{1}{3}$. This estimate is uniform in $\varepsilon$ by Assumption 1.1 as well as (E1), (E3) and (E4). Similarly,
$\left|\int_{0}^{t}\int_{\mathbb{R}\setminus[-M,M]}u\,\partial_{x}s^{\alpha}\,\partial_{x}\mathcal{V}^{1}[u,v]\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau\right|\leq\frac{C}{M^{\kappa}}$
for some constant $C$. This proves the convergence (5.15) by taking the limit $M\to\infty$. Analogously,
@@
It follows that $\limsup_{\varepsilon\to 0}\int_{0}^{t}\int_{\mathbb{R}}|\partial_{x}s_{\varepsilon}^{\alpha}-\partial_{x}s^{\alpha}|^{2}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau\leq 0$ for a.e. $t\in(0,T)$. We conclude since $T>0$ is arbitrary and the proof of the theorem is complete.
@@
One may ask whether the strong convergence of $\{u_{\varepsilon}\}$ and $\{v_{\varepsilon}\}$ can be extended to the whole domain $0,T]\times\mathbb{R}$. A natural approach would be to exploit the energy $(u,v)\mapsto\int_{\mathbb{R}}(u\log u+v\log v)\mathop{}\!\mathrm{d}x$, in analogy with the use of the dissipation of $s\mapsto\int_{\mathbb{R}}s\log s\mathop{}\!\mathrm{d}x$ to prove the strong convergence of $\{\partial_{x}s_{\varepsilon}^{\alpha}\}$. The main difficulty is that, although all the terms in the PDE for the limit $u$ have been identified, it remains unclear how to justify rigorously testing this PDE by $\log u$ (and similarly testing the equation for $v$ by $\log v$), since $u$ and $v$ possess only $L^{p}((0,T)\times\mathbb{R})$ regularity. This argument could also show strong convergence of $\{\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\}$.
@@
$\begin{split}\frac{1}{3}\,\overline{\,\frac{u^{4}}{s^{2}}\,}\,\partial_{x}(\mathcal{V}^{1}[u,v]-&\mathcal{V}^{2}[u,v])=\\ &=\frac{1}{\alpha}\,\overline{\frac{u}{s}\partial_{x}s^{\alpha}}\,\overline{\,\frac{u^{2}}{s}\,}-\frac{1}{\alpha}\,u\,\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}+\frac{1}{3}\,u\,\overline{\,\frac{u^{3}}{s^{2}}\,}\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]),\end{split}$ (5.18)
@@
where $\overline{\,\frac{u^{4}}{s^{2}}\,}$, $\overline{\frac{u}{s}\partial_{x}s^{\alpha}}$, $\overline{\,\frac{u^{2}}{s}\,}$, $\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}$, $\overline{\,\frac{u^{3}}{s^{2}}\,}$ are the weak limits defined in (C13) and (C14). We apply Theorem 4.3 with
$U_{\varepsilon}:=u_{\varepsilon},$ $F_{\varepsilon}:=\frac{1}{\alpha}\,\frac{u_{\varepsilon}}{s_{\varepsilon}}\partial_{x}s_{\varepsilon}^{\alpha}+u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}],$
$V_{\varepsilon}:=\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}},$ $G_{\varepsilon}:=\frac{1}{\alpha}\,\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}^{\alpha}+\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\frac{u_{\varepsilon}^{3}}{3\,s_{\varepsilon}^{2}}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]),$
$R_{\varepsilon}:=\varepsilon\,\partial^{2}_{x}u_{\varepsilon},$ $S^{1}_{\varepsilon}:=h_{\varepsilon},\quad S^{2}_{\varepsilon}:=g_{\varepsilon},$
where $g_{\varepsilon}$, $h_{\varepsilon}$ are defined in Proposition 3.7. This is allowed because
  • $\{U_{\varepsilon}\}$, $\{V_{\varepsilon}\}$ are bounded in $L^{\infty}(0,T;L^{2}(\mathbb{R}))$ by (E4),
  • $\{F_{\varepsilon}\}$, $\{G_{\varepsilon}\}$ are bounded in $L^{2}((0,T)\times\mathbb{R})$ thanks to (E3), (E4) and Assumption 1.1 which implies that $\{\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\}$, $\{\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]\}$ are bounded in $L^{\infty}((0,T)\times\mathbb{R})$,
  • $\{R_{\varepsilon}\}$ is strongly compact in $L^{2}(0,T;H^{-1}(\mathbb{R}))$ since $R_{\varepsilon}:=\partial_{x}(\varepsilon\,\partial_{x}u_{\varepsilon})$ with $\varepsilon\,\partial_{x}u_{\varepsilon}\!\to\!0$ strongly in $L^{2}((0,T)\times\mathbb{R})$ by (E8),
  • $\{S^{1}_{\varepsilon}\}$ is strongly compact in $L^{2}(0,T;H^{-1}(\mathbb{R}))$ by Proposition 3.7,
  • $\{S^{2}_{\varepsilon}\}$ is bounded in $\mathcal{M}_{\operatorname{loc}}([0,T]\times\mathbb{R})$ by Proposition 3.7.
By Theorem 4.3, the weak limit of $F_{\varepsilon}\,V_{\varepsilon}-U_{\varepsilon}\,G_{\varepsilon}$ is $\overline{F}\,\,\overline{V}-\overline{U}\,\,\overline{V}$. We compute
@@
$F_{\varepsilon}\,V_{\varepsilon}-U_{\varepsilon}\,G_{\varepsilon}=$ $\left(\frac{1}{\alpha}\,\frac{u_{\varepsilon}}{s_{\varepsilon}}\partial_{x}s_{\varepsilon}^{\alpha}+u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\right)\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}$
$-u_{\varepsilon}\,\left(\frac{1}{\alpha}\,\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}^{\alpha}+\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\frac{u_{\varepsilon}^{3}}{3\,s_{\varepsilon}^{2}}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])\right)$
$=$ $\frac{u_{\varepsilon}^{4}}{3\,s_{\varepsilon}^{2}}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]).$
@@
$F_{\varepsilon}\,V_{\varepsilon}-U_{\varepsilon}\,G_{\varepsilon}\rightharpoonup\frac{1}{3}\,\overline{\,\frac{u^{4}}{s^{2}}\,}\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]).$
Moreover, by (C2), (C5), (C13), (C14)
$\overline{F}\,\,\overline{V}=\left(\frac{1}{\alpha}\,\overline{\frac{u}{s}\partial_{x}s^{\alpha}}+u\,\partial_{x}\mathcal{V}^{1}[u,v]\right)\,\overline{\,\frac{u^{2}}{s}\,}$,
$\overline{U}\,\,\overline{G}=u\,\left(\frac{1}{\alpha}\,\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}+\overline{\,\frac{u^{2}}{s}\,}\partial_{x}\mathcal{V}^{1}[u,v]-\frac{1}{3}\overline{\,\frac{u^{3}}{s^{2}}\,}\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v])\right),$
@@
$\overline{F}\,\,\overline{V}-\overline{U}\,\,\overline{G}=\frac{1}{\alpha}\,\overline{\frac{u}{s}\partial_{x}s^{\alpha}}\,\overline{\,\frac{u^{2}}{s}\,}-\frac{1}{\alpha}\,u\,\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}+\frac{1}{3}\,u\,\overline{\,\frac{u^{3}}{s^{2}}\,}\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]).$
@@
$\frac{s^{2}}{3}\,\overline{\,\frac{u^{4}}{s^{4}}\,}\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v])=\\ =\frac{s}{\alpha}\,\overline{\frac{u}{s}\partial_{x}s^{\alpha}}\,\overline{\,\frac{u^{2}}{s^{2}}\,}-\frac{s}{\alpha}\,\frac{u}{s}\,\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}+\frac{s^{2}}{3}\,\frac{u}{s}\,\overline{\,\frac{u^{3}}{s^{3}}\,}\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]).$
@@
$q_{\varepsilon}^{3}=((q_{\varepsilon}-q)+q)^{3}=(q_{\varepsilon}-q)^{3}+3\,(q_{\varepsilon}-q)^{2}\,q+3\,(q_{\varepsilon}-q)\,q^{2}+q^{3}$

### 16 Apr 2026 v2

@@ 5.1. The support of $\pi_{t,x}$ is (almost always) at most a line
Indeed, we split $\mathbb{R}$ for [R,R][-R,R] and $\mathbb{R}\setminus[-R,R]$. For all $R>0$, we have covergence of the integral over [R,R][-R,R] thanks to (C1), (C5) and because $\overline{\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}s_{\varepsilon}^{\alpha}}=\frac{u}{s}\,\partial_{x}s^{\alpha}$. On $\mathbb{R}\setminus[-R,R]$ we use that the integral is small
$\left|\int_{0}^{t}\int_{\mathbb{R}\setminus[-R,R]}u_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau\right|\leq\|\partial_{x}\mathcal{V}^{1}\|_{L^{\infty}}\,\|\partial_{x}s_{\varepsilon}^{\alpha}\|_{L^{2}((0,T)\times\mathbb{R})}\,\frac{\|u_{\varepsilon}\,|x|^{\frac{1}{2}}\|_{L^{2}((0,T)\times\mathbb{R})}}{R^{\frac{1}{2}}}.$
This estimate is uniform in $\varepsilon$ by Assumption 1.1 as well as (E1), (E3) and (E4). Similarly,
$\left|\int_{0}^{t}\int_{\mathbb{R}\setminus[-R,R]}u\,\partial_{x}s^{\alpha}\,\partial_{x}\mathcal{V}^{1}[u,v]\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau\right|\leq\frac{C}{R^{\frac{1}{2}}}$
for some constant $C$. This proves the convergence (5.15) by taking the limit $R\to\infty$. Analogously,
@@
It follows that $\limsup_{\varepsilon\to 0}\int_{0}^{t}\int_{\mathbb{R}}|\partial_{x}s_{\varepsilon}^{\alpha}-\partial_{x}s^{\alpha}|^{2}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau\leq 0$ for a.e. $t\in(0,T)$. We conclude since $T>0$ is arbitrary.
We now prove the strong convergence of $\{u_{\varepsilon}\}$, $\{v_{\varepsilon}\}$ in $L^{p}(\mathcal{S})$ and a.e. on $\mathcal{S}$, where
$\mathcal{S}:=[0,T]~\times~\mathbb{R}\setminus\{(t,x):s(t,x)>0\mbox{ and }\partial_{x}\mathcal{V}^{1}[u,v]=\partial_{x}\mathcal{V}^{2}[u,v]\}.$
Let $\{\mu_{t,x}\}_{t,x}$ be the Young measure of $\{u_{\varepsilon}\}$, i.e. whenever $f(t,x,\lambda):[0,T]\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function (measurable in $(t,x)$ and continuous in $\lambda$) and $\{f(t,x,u_{\varepsilon})\}$ is weakly compact in $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$, $f(t,x,u_{\varepsilon})\rightharpoonup\int_{[0,\infty)}f(t,x,\lambda)\mathop{}\!\mathrm{d}\mu_{t,x}\mbox{ weakly in }L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R}),$ see [46, Chapter 6]. We will prove that for $(t,x)\in\mathcal{S}$ we have $\mu_{t,x}=\delta_{u(t,x)}$, which implies the claim since $|u_{\varepsilon}-u|^{p}\rightharpoonup\int_{[0,\infty)}|\lambda-u(t,x)|^{p}\mathop{}\!\mathrm{d}\mu_{t,x}(\lambda)=0$ for $p\geq 1$. We consider two cases.
  • When $s(t,x)>0$ and $\partial_{x}\mathcal{V}^{1}[u,v]\neq\partial_{x}\mathcal{V}^{2}[u,v]$, by Propositions 5.1 and 5.2 we have $\pi_{t,x}((\mathbb{R}\setminus\{0\})\times\mathbb{R})=0$, so that $\lambda_{1}=0$ for a.e.-$\pi_{t,x}$ $(\lambda_{1},\lambda_{2})$. In particular, by (C13) and $\overline{\,\frac{u^{2}}{s^{2}}\,}-\frac{u^{2}}{s^{2}}=\int_{\mathbb{R}^{2}}\lambda_{1}^{2}\mathop{}\!\mathrm{d}\pi_{t,x}(\lambda_{1},\lambda_{2})$,
@@
One may ask whether the strong convergence of $\{u_{\varepsilon}\}$ and $\{v_{\varepsilon}\}$ can be extended to the whole domain $0,T]\times\mathbb{R}$. The compensated compactness approach fails at points where $\partial_{x}\mathcal{V}^{1}[u,v]=\partial_{x}\mathcal{V}^{2}[u,v]$ (which implies $\mathcal{F}=0$ with $\mathcal{F}$ defined in (5.1)), since (5.8) reduces to the trivial identity $0=0$ once $\{\partial_{x}s_{\varepsilon}^{\alpha}\}$ is known to converge strongly. Similarly, in the alternative argument of Section 5.4, (5.18) degenerates to the same trivial identity. A natural approach would be to exploit the energy $(u,v)\mapsto\int_{\mathbb{R}}(u\log u+v\log v)\mathop{}\!\mathrm{d}x$, in analogy with the use of the dissipation of $s\mapsto\int_{\mathbb{R}}s\log s\mathop{}\!\mathrm{d}x$ to prove the strong convergence of $\{\partial_{x}s_{\varepsilon}^{\alpha}\}$. The main difficulty is that, although all the terms in the PDE for the limit $u$ have been identified, it remains unclear how to justify rigorously testing this PDE by $\log u$ (and similarly testing the equation for $v$ by $\log v$), since $u$ and $v$ possess only $L^{p}((0,T)\times\mathbb{R})$ regularity. This argument could also show strong convergence of $\{\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}\}$.
@@
$\begin{split}\frac{1}{3}\,\overline{\,\frac{u^{4}}{s^{2}}\,}\,\partial_{x}(\mathcal{V}^{1}[u,v]-&\mathcal{V}^{2}[u,v])\\ &=\frac{1}{\alpha}\,\overline{\frac{u}{s}\partial_{x}s^{\alpha}}\,\overline{\,\frac{u^{2}}{s}\,}-\frac{1}{\alpha}\,u\,\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}+\frac{1}{3}\,u\,\overline{\,\frac{u^{3}}{s^{2}}\,}\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]),\end{split}$ (5.18)
@@
where $\overline{\,\frac{u^{4}}{s^{2}}\,}$, $\overline{\frac{u}{s}\partial_{x}s^{\alpha}}$, $\overline{\,\frac{u^{2}}{s}\,}$, $\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}$, $\overline{\,\frac{u^{3}}{s^{2}}\,}$ are the weak limits defined in (C13) and (C14). As in Section 5.2, we apply Theorem 4.3 to the vector fields
${\bf a_{\varepsilon}}$ $=\left(u_{\varepsilon},-\frac{1}{\alpha}\,\frac{u_{\varepsilon}}{s_{\varepsilon}}\partial_{x}s_{\varepsilon}^{\alpha}-u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\right),$
${\bf b_{\varepsilon}}$ $=\left(\frac{1}{\alpha}\,\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}^{\alpha}+\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\frac{u_{\varepsilon}^{3}}{3\,s_{\varepsilon}^{2}}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]),\,\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\right).$
Let us explain why the assumptions of Theorem 4.3 are satisfied. Clearly, both $\{\bf a_{\varepsilon}\}$ and $\{\bf b_{\varepsilon}\}$ converge weakly in $L^{2}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ to some $\bf a$} and $\bf b$} by (C1), (C2), (C5), (C13), and (C14) in Proposition 3.4. Moreover, $\operatorname{div}{\bf a_{\varepsilon}}=\varepsilon\,\partial^{2}_{x}u_{\varepsilon}$ converges strongly to 0 in $H^{-1}((0,T)\times\mathbb{R})$ because, by (E8), $\varepsilon\,\partial_{x}u_{\varepsilon},\varepsilon\,\partial_{x}v_{\varepsilon}\to 0$ strongly in $L^{2}((0,T)\times\mathbb{R})$. Regarding $\operatorname{curl}{\bf b_{\varepsilon}}$, we have $\operatorname{curl}{\bf b_{\varepsilon}}=-g_{\varepsilon}-h_{\varepsilon}$, where $\{g_{\varepsilon}\}$ and $\{h_{\varepsilon}\}$ are defined in Proposition 3.7. We know that $\{g_{\varepsilon}\}$ is bounded in $L^{1}_{\operatorname{loc}}([0,T]\times\mathbb{R})$, which compactly embeds into $W^{-1,p}_{\operatorname{loc}}([0,T]\times\mathbb{R})$ for $p<2$ [29, Theorem 6, p. 7], while $\{h_{\varepsilon}\}$ is strongly compact in $H^{-1}((0,T)\times\mathbb{R})$. Combining these two facts, $\{\operatorname{curl}{\bf b_{\varepsilon}}\}$ is strongly compact in $(W^{1,\infty}_{0}(\Omega))^{*}$ for each bounded $\Omega\subset[0,T]\times\mathbb{R}$. Finally, the product $\{{\bf a_{\varepsilon}}\cdot{\bf b_{\varepsilon}}\}$ is uniformly integrable in $L^{1}(\Omega)$ for $\Omega$ as above. Indeed, writing ${\bf a_{\varepsilon}}=(a^{1}_{\varepsilon},a^{2}_{\varepsilon})$ and ${\bf b_{\varepsilon}}=(b^{1}_{\varepsilon},b^{2}_{\varepsilon})$, $\{a^{1}_{\varepsilon}\}$ and $\{b^{2}_{\varepsilon}\}$ are bounded in $L^{\infty}((0,T)\times\mathbb{R})$ by (E4), while $\{a^{2}_{\varepsilon}\}$ and $\{b^{1}_{\varepsilon}\}$ are bounded at least in $L^{2}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ by (E3), (E4), and Assumption 1.1. It follows that $\{{\bf a_{\varepsilon}}\cdot{\bf b_{\varepsilon}}\}$ is bounded in $L^{2}(\Omega)$, which implies uniform integrability in $L^{1}(\Omega)$.
We compute
@@
${\bf a_{\varepsilon}}\cdot{\bf b_{\varepsilon}}=$ $\,u_{\varepsilon}\left(\frac{1}{\alpha}\,\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}^{\alpha}+\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\frac{u_{\varepsilon}^{3}}{3\,s_{\varepsilon}^{2}}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])\right)$
$-\left(\frac{1}{\alpha}\,\frac{u_{\varepsilon}}{s_{\varepsilon}}\partial_{x}s_{\varepsilon}^{\alpha}+u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\right)\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}$
$=$ $\,-\frac{u_{\varepsilon}^{4}}{3\,s_{\varepsilon}^{2}}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]).$
@@
${\bf a_{\varepsilon}}\cdot{\bf b_{\varepsilon}}\rightharpoonup-\frac{1}{3}\,\overline{\,\frac{u^{4}}{s^{2}}\,}\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]).$
Moreover, by (C2), (C5), (C13), (C14), we can identify also products of the weak limits ${\bf a}=(a^{1},a^{2})$ and ${\bf b}=(b^{1},b^{2})$ as
$a^{1}\,b^{1}=u\,\left(\frac{1}{\alpha}\,\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}+\overline{\,\frac{u^{2}}{s}\,}\partial_{x}\mathcal{V}^{1}[u,v]-\frac{1}{3}\overline{\,\frac{u^{3}}{s^{2}}\,}\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v])\right),$
$a^{2}\,b^{2}=-\left(\frac{1}{\alpha}\,\overline{\frac{u}{s}\partial_{x}s^{\alpha}}+u\,\partial_{x}\mathcal{V}^{1}[u,v]\right)\,\overline{\,\frac{u^{2}}{s}\,}$,
@@
${\bf a}\cdot{\bf b}=-\frac{1}{\alpha}\,\overline{\frac{u}{s}\partial_{x}s^{\alpha}}\,\overline{\,\frac{u^{2}}{s}\,}+\frac{1}{\alpha}\,u\,\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}-\frac{1}{3}\,u\,\overline{\,\frac{u^{3}}{s^{2}}\,}\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]).$
@@
$\frac{s^{2}}{3}\,\overline{\,\frac{u^{4}}{s^{4}}\,}\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v])\\ =\frac{s}{\alpha}\,\overline{\frac{u}{s}\partial_{x}s^{\alpha}}\,\overline{\,\frac{u^{2}}{s^{2}}\,}-\frac{s}{\alpha}\,\frac{u}{s}\,\overline{\frac{u^{2}}{s^{2}}\,\partial_{x}s^{\alpha}}+\frac{s^{2}}{3}\,\frac{u}{s}\,\overline{\,\frac{u^{3}}{s^{3}}\,}\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]).$
@@
$q_{\varepsilon}^{3}=((q_{\varepsilon}-q)+q)^{3}=(q_{\varepsilon}-q)^{3}+3\,(q_{\varepsilon}-q)^{2}\,q+3\,(q_{\varepsilon}-q)\,q^{2}+q^{3},$